OpenStudy (anonymous):

x = 3y - 4 2x - y = 7 How I do I solve this using the substitution method?

OpenStudy (campbell_st):

take the 1st equation and substitiute it into the 2nd equation and you will have 2(3y - 4) - y = 7 distribute and you will have 6y - 8 - y = 7 or 5y - 8 = 7 solve for y when you get y, substitute into either equation to get the value of x.